# Roton-mediated soliton bound states in binary dipolar condensates

### v1

@@ I Introduction
A parallel research thrust in multicomponent BECs is the study of vector solitons, excitations that span multiple components through intercomponent coupling. In two-component (binary) mixtures, the excitation spectrum splits into a density branch, corresponding to in-phase density fluctuations, and a spin branch, corresponding to out-of-phase fluctuations. The spin branch, characterized by fluctuations of the spin density $n_{1}-n_{2}$, plays a central role in both the internal structure of vector solitons and their interactions. An important experimentally realized example in binary condensates is the dark-antidark soliton [28, 29, 30, 31, 32, 33, 34], a solitary wave in which one miscible component exhibits a phase step and density dip, while the other shows a corresponding density peak. Of particular interest are bound states of oppositely polarized dark-antidark soliton pairs (also termed dark-dark solitons) [35, 36, 37], which have been observed experimentally in counter-flowing BECs [28, 29]. These bound states result from attractive inter-soliton interactions, and give rise to dynamics and stability properties distinct from conventional single-component solitons [38]. Setting the stage for the present work, we emphasize that although both binary condensates and single-component dipolar condensates can host solitary wave bound states, the underlying physical mechanisms differ. Moreover, dipolar mixtures have recently been cooled to quantum degeneracy [39, 40, 41, 42, 43, 44, 45], and their capacity to realize vector solitons remains unexplored.
@@ II.1 System and interactions
with $g_{ij}^{\rm dd}=4\pi a_{ij}^{\rm dd}\hbar^{2}/m$, where $a_{ij}^{\rm dd}=\mu_{0}\mu_{i}\mu_{j}m/12\pi\hbar$ are the dipole lengths associated with magnetic moments $\mu_{i}$. For simplicity, we assume both components have the same mass $m$. The quantity $\vartheta$ represents the angle between the dipole alignment direction $\hat{z}$ and the vector $\mathbf{r}=\mathbf{x}-\mathbf{x}^{\prime}$, representing the relative position of the interacting particles.
@@ III.1 Isolated dark-antidark solitons
Figure 1 displays the (a) density and (b) phase distributions of a dark-antidark soliton at rest. The dipolar component (component 1) hosts a dark soliton with a $\pi$ phase step, while the nondipolar component (component 2) forms an antidark structure with uniform phase. At the soliton core, the spin density $s(x)$ becomes negative [$s(x)<0$)], which we define as a negative-core dark-antidark soliton. Conversely, when the soliton core exhibits $s(x)>0$, we refer to it as a positive-core dark-antidark soliton. Figure 1(c) shows that while this excitation generates small modulations in the total density, Fig. 1(d) reveals pronounced oscillations of the spin density. The characteristic spacing between the resulting peaks and troughs can be understood in terms of the system’s dispersion relations.
@@ III.2 Derivation of dispersion relations
Binary condensates exhibit two gapless Goldstone dispersion relations, arising from the two spontaneously broken gauge symmetries associated with each order parameter. Because of the intercomponent coupling, one branch is predominantly associated with fluctuations of the total density $n(x)=n_{1}(x)+n_{2}(x)$, while the other corresponds to fluctuations of the spin density $s(x)=n_{1}(x)-n_{2}(x)$. Using Bogoliubov theory, we derive an expression for the density (++) and spin (-) dispersion relation branches (see Appendix A):
$\epsilon_{\pm}^{2}(k)=\frac{\epsilon_{1}^{2}+\epsilon_{2}^{2}}{2}\pm\frac{1}{2}\sqrt{(\epsilon_{1}^{2}-\epsilon_{2}^{2})^{2}+4G_{12}G_{12}\frac{\hbar^{4}k^{4}}{m_{1}m_{2}}},$ (5)
@@
$\epsilon_{i}^{2}(k)=\frac{\hbar^{2}k^{2}}{2m_{i}}\left(\frac{\hbar^{2}k^{2}}{2m_{i}}+2G_{ii}(k)\right),$ (6)
@@ III.3 Role of dispersion relations in soliton structure
For nondipolar mixtures [Fig. 2(a1)], the spin branch smoothly approaches the free-particle dispersion (thin dashed line), with the spin healing length $\xi_{s}=\hbar/\sqrt{2mn_{0}\delta g/\pi l^{2}}$ diverging at the transition where $\delta g\equiv(\sqrt{g_{11}g_{22}}-g_{12})=0$. In contrast, the dipolar mixture [Fig. 2(b1)] develops a spin-roton minimum [49, 50, 51, 59], which deepens with increasing $a_{12}$ and bends the branch below the free-particle curve, resulting in effectively repulsive interactions at small $k$ but attractive interactions at large $k$. In both systems, the density branches $\epsilon_{+}$ (dashed lines) depend only weakly on $a_{12}$.
The corresponding spin-density soliton profiles reflect these contrasting dispersion relations  [^1] Note that while the dispersion relation is calculated for the homogeneous system, the numerical simulations introduce a small finite-size effect due to the presence of the soliton. As a result, the critical value of $a_{12}$ is approximately 1% higher in Figs. 2(a2,b2) that in Figs. 2(a1,b1). In the nondipolar case [Fig. 2(a2)], the soliton width grows with increasing $\xi_{s}$, while the central spin density approaches $s=-2n_{0}$, consistent with a magnetic soliton of nearly constant total density [61]. In dipolar mixtures [Fig. 2(b2)], the roton generates spin-density oscillations of wavelength $\lambda_{\text{rot}}/l\approx 2.7$ around the soliton core. The width of the central minimum remains $\sim\lambda_{\text{rot}}$, but the decay length of the modulations increases with increasing roton depth. Note that at the miscible-immiscible threshold, spin-density fluctuations cost zero energy, signally a transition to the alternating-domain supersolid phase [62, 63, 59]. The presence of spin rotons not only shapes individual solitons but also plays a crucial role in mediating the interactions between soliton pairs, setting the stage for the formation of multiple bound states.
@@ III.4 Multiple bound states and soliton-pair interactions
The upper panels of Fig. 3 illustrate representative examples of three types of excited bound states with distinct equilibrium separations. Panel (i) shows solitons oscillating through one another at $a_{12}/a_{0}=95$, analogous to magnetic soliton bound states in binary nondipolar condensates [38]. Panels (ii) and (iii) depict long-range bound states oscillating about equilibrium positions $\Delta x/l\approx 2.7$ and $\Delta x/l\approx 5.4$ for $a_{12}/a_{0}=70$ and $a_{12}/a_{0}=95$, respectively, where $\Delta x=x_{2}-x_{1}$ with $x_{i}$ the position of soliton $i$. Because dipolar interactions break integrability, these structures are not true solitons in the strict sense and radiate energy through phonon emission. Dissipation-driven acceleration eventually pushes oscillating solitons beyond their escape velocity [26, 64], as seen in Fig. 3(i) at $t\omega\approx 60$. This counterintuitive behavior arises from the negative effective mass of dark-antidark solitons.
@@
where $E_{\text{tot}}$ is the system energy of two zero-velocity solitons separated by a distance $\Delta x$, and $E_{\text{sep}}$ is the system energy of two well-separated solitons, ensuring $E=0$ at maximal separation. For dark-antidark solitons, the effective mass satisfies $m_{\rm{sol}}<0$ [61], and its sign is included in Eq. (8) to render the direction of the restoring force for bound-state oscillations in Fig. 3 more transparent.
@@
$\text{slog}(x)=\text{sign}(x)\log\left(\sqrt{x^{2}+1}\right),$ (10)
@@ III.5 Origin of multiple bound states
Near the roton instability, $a_{12}^{\text{c}}/a_{0}\approx 110$, the decay length of the spatial spin density oscillations around the soliton core grows, so too does the number of alternating peaks and troughs of the inter-soliton potential energy. For soliton pairs with like sign, potential minima arise at $|\Delta x|\approx(I+1/2)\lambda_{\text{rot}}$, whereas for opposite-sign pairs they occur at $|\Delta x|\approx I\lambda_{\text{rot}}$, with $I=0,1,2,\dots$. In this work we restrict our attention to the first three such minima.
@@ III.6 Soliton collisions
True solitons can, in principle, emerge unscathed from collisions. On the one hand, the presence of dipolar interactions breaks integrability, so perfectly elastic collisions are not expected. On the other hand, collisions between dark-antidark solitons may still be nearly elastic and thus serve as a probe of the inter-soliton potential.
Figure 5 shows four dark-antidark solitons in each panel—two slowly moving negative-core solitons and two stationary positive-core solitons—in (a) the nondipolar reference system and (b) a binary dipolar condensate. In the nondipolar case, like-sign solitons bounce off each other, whereas opposite-sign solitons pass through (as predicted in Ref. [38]). With dipolar interactions present, however, they always bounce, independent of relative sign. The bouncing arises from maxima in the inter-soliton potential $V$, which act as effective barriers. For like-sign solitons, these maxima occur at integer multiples of $\lambda_{\text{rot}}$; for opposite-sign solitons, half-integer multiples (see Fig. 4). Consequently, for the collisions shown in Fig. 5(b), the minimum separation before bouncing differs: $\Delta x\gtrsim\lambda_{\text{rot}}$ for like-sign collisions and $\Delta x\gtrsim 1.5\lambda_{\text{rot}}$ for opposite-sign collisions, suggesting that in both cases the second maxima are relevant at these velocities. These distinct collisional properties—especially when contrasted with the nondipolar case—could provide direct experimental evidence of the spin roton.
@@ IV Conclusion
Acknowledgements:— The authors acknowledge T. Billam, T. Bland, F. Ferlaino, L. B. Giacomelli, A. Madhusudan, N. Masalaeva, N. Parker, E. Poli, C. Qu and P. Senarath Yapa for insightful discussions. This research was funded by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/P 36850-N].
Appendix A Bogoliubov theory
@@
$\epsilon_{\pm}^{2}(k)=\frac{\epsilon_{1}^{2}+\epsilon_{2}^{2}}{2}\pm\frac{1}{2}\sqrt{(\epsilon_{1}^{2}-\epsilon_{2}^{2})^{2}+4G_{12}G_{12}\frac{\hbar^{4}k^{4}}{m_{1}m_{2}}},$ (14)
@@
$\epsilon_{i}^{2}(k)=\frac{\hbar^{2}k^{2}}{2m_{i}}\left(\frac{\hbar^{2}k^{2}}{2m_{i}}+2G_{ii}(k)\right),$ (15)

### v2

@@ I Introduction
A parallel research thrust in multicomponent BECs is the study of vector solitons, excitations that span multiple components through intercomponent coupling. In two-component (binary) mixtures, the excitation spectrum splits into a density branch, corresponding to in-phase density fluctuations, and a spin branch, corresponding to out-of-phase fluctuations. The spin branch, characterized by fluctuations of the spin density $n_{1}-n_{2}$, plays a central role in both the internal structure of vector solitons and their interactions. An important experimentally realized example in binary condensates is the dark-antidark soliton [28, 29, 30, 31, 32, 33, 34], a solitary wave in which one miscible component exhibits a phase step and density dip, while the other shows a corresponding density peak. Of particular interest are bound states of oppositely polarized dark-antidark soliton pairs (also termed dark-dark solitons) [35, 36, 37], which have been observed experimentally in counter-flowing BECs [28, 29]. These bound states result from attractive inter-soliton interactions, and give rise to dynamics and stability properties distinct from conventional single-component solitons [38]. Setting the stage for the present work, we emphasize that although both binary condensates and single-component dipolar condensates can host solitary wave bound states, the underlying physical mechanisms differ. Moreover, dipolar mixtures have recently been cooled to quantum degeneracy [39, 40, 41, 42, 43, 44, 45, 46, 47], and their capacity to realize vector solitons remains unexplored.
@@ II.1 System and interactions
with $g_{ij}^{\rm dd}=4\pi a_{ij}^{\rm dd}\hbar^{2}/m$, where $a_{ij}^{\rm dd}=\mu_{0}\mu_{i}\mu_{j}m/12\pi\hbar^{2}$ are the dipole lengths associated with magnetic moments $\mu_{i}$. For simplicity, we assume both components have the same mass $m$. The quantity $\vartheta$ represents the angle between the dipole alignment direction $\hat{z}$ and the vector $\mathbf{r}=\mathbf{x}-\mathbf{x}^{\prime}$, representing the relative position of the interacting particles.
@@ III.1 Isolated dark-antidark solitons
Figure 1 displays the (a) density and (b) phase distributions of a dark-antidark soliton at rest. The dipolar component (component 1) hosts a dark soliton with a $\pi$ phase step, while the nondipolar component (component 2) forms an antidark structure with uniform phase. Figure 1(c) shows that, while this excitation generates small modulations in the total density $n(x)=n_{1}(x)+n_{2}(x)$, Fig. 1(d) reveals pronounced oscillations of the spin density $s(x)=n_{1}(x)-n_{2}(x)$. Since the sign of the spin density also oscillates, we define the soliton sign as the sign of $s(x)$ at its core, determined by which components carries the phase jump. For the present example, at the core $s(x)<0$, which we define as a negative-core dark-antidark soliton. Conversely, when the soliton core exhibits $s(x)>0$, we refer to it as a positive-core dark-antidark soliton. The characteristic spacing between the resulting peaks and troughs of $s(x)$ can be understood in terms of the system’s dispersion relations.
@@ III.2 Derivation of dispersion relations
Binary condensates exhibit two gapless Goldstone dispersion relations, arising from the two spontaneously broken gauge symmetries associated with each order parameter. Because of the intercomponent coupling, one branch is predominantly associated with fluctuations of the total density $n(x)$, while the other corresponds to fluctuations of the spin density $s(x)$. Using Bogoliubov theory, we derive an expression for the density (++) and spin (-) dispersion relation branches (see Appendix B):
$\epsilon_{\pm}^{2}(k)=\frac{\epsilon_{1}^{2}+\epsilon_{2}^{2}}{2}\pm\frac{1}{2}\sqrt{(\epsilon_{1}^{2}-\epsilon_{2}^{2})^{2}+4G_{12}G_{12}\frac{\hbar^{4}k^{4}}{m^{2}}},$ (5)
@@
$\epsilon_{i}^{2}(k)=\frac{\hbar^{2}k^{2}}{2m}\left(\frac{\hbar^{2}k^{2}}{2m}+2G_{ii}(k)\right),$ (6)
@@ III.3 Role of dispersion relations in soliton structure
For nondipolar mixtures [Fig. 2(a1)], the spin branch smoothly approaches the free-particle dispersion (thin dashed line), with the spin healing length $\xi_{s}=\hbar/\sqrt{2mn_{0}\delta g/\pi l^{2}}$ diverging at the transition where $\delta g\equiv(\sqrt{g_{11}g_{22}}-g_{12})=0$. In contrast, the dipolar mixture [Fig. 2(b1)] develops a spin-roton minimum [51, 52, 53, 62], which deepens with increasing $a_{12}$ and bends the branch below the free-particle (non-interacting) curve, corresponding to effectively repulsive interactions at small $k$ but attractive interactions at large $k$. This can be understood by noting from Eqs. (5)-(7) that deviations from the free particle energy, $\epsilon_{0}=\hbar^{2}k^{2}/2m$, arise from $G_{ij}(k)$, and the changing sign of $\epsilon_{-}(k)-\epsilon_{0}(k)$ is caused by the momentum dependence of the dipole-dipole interactions $U_{ij}(k)$  [^2] For discussions of the momentum-dependent effective interaction in single-component dipolar gases, see, e.g., Refs. [14, 70, 71]. . In both systems, the density branches $\epsilon_{+}$ (dashed lines) depend only weakly on $a_{12}$.
The corresponding spin-density soliton profiles reflect these contrasting dispersion relations  [^3] Note that while the dispersion relation is calculated for the homogeneous system, the numerical simulations introduce a small finite-size effect due to the presence of the soliton. As a result, the critical value of $a_{12}$ is approximately 1% higher in Figs. 2(a2,b2) that in Figs. 2(a1,b1). In the nondipolar case [Fig. 2(a2)], the soliton width grows with increasing $\xi_{s}$, while the central spin density approaches $s=-2n_{0}$, consistent with a magnetic soliton of nearly constant total density [65]. In dipolar mixtures [Fig. 2(b2)], the roton generates spin-density oscillations of wavelength $\lambda_{\text{rot}}/l\approx 2.7$ around the soliton core. The width of the central minimum remains $\sim\lambda_{\text{rot}}$, but the decay length of the modulations increases with increasing roton depth. Note that at the miscible-immiscible threshold, spin-density fluctuations cost zero energy, signally a transition to the alternating-domain supersolid phase [66, 67, 62]. The presence of spin rotons not only shapes individual solitons but also plays a crucial role in mediating the interactions between two solitons, setting the stage for the formation of multiple bound states.
@@ III.4 Multiple bound states and soliton-pair interactions
The upper panels of Fig. 3 illustrate representative examples of three types of excited bound states with distinct equilibrium separations. Panel (i) shows solitons oscillating through one another at $a_{12}/a_{0}=95$, analogous to magnetic soliton bound states in binary nondipolar condensates [38]. Panels (ii) and (iii) depict long-range bound states oscillating about equilibrium positions $\Delta x/l\approx 2.7$ and $\Delta x/l\approx 5.4$ for $a_{12}/a_{0}=70$ and $a_{12}/a_{0}=95$, respectively, where $\Delta x=x_{2}-x_{1}$ with $x_{i}$ the position of soliton $i$. Because dipolar interactions break integrability, these structures are not true solitons in the strict sense and radiate energy through phonon emission [26]. Such phonons can be seen as fluctuations in the spin density far from the solitons in Fig. 3(i), even for early times. Dissipation-driven acceleration eventually pushes oscillating solitons beyond their escape velocity [26, 68], as seen in Fig. 3(i) at $t\omega\approx 60$. This counterintuitive behavior indicates that dark-antidark solitons in binary dipolar BECs have negative effective masses, similar to dark solitons in single-component dipolar BECs [26, 27] and dark-antidark soltions in binary BECs [65]. We confirm this numerically for our system in Appendix C.
@@
where $E_{\text{tot}}$ is the system energy of two zero-velocity solitons separated by a distance $\Delta x$, and $E_{\text{sep}}$ is the system energy of two well-separated solitons, ensuring $E=0$ at maximal separation. For dark-antidark solitons, the effective mass satisfies $m_{\rm{sol}}<0$, and its sign is included in Eq. (8) to render the direction of the restoring force for bound-state oscillations in Fig. 3 more transparent.
@@
$\text{slog}(x)=\text{sgn}(x)\log_{10}\left(\sqrt{x^{2}+1}\right),$ (10)
@@ III.5 Origin of multiple bound states
When approaching the roton instability, $a_{12}^{\text{c}}/a_{0}\approx 110$, the decay length of the spatial spin density oscillations around the soliton core grows, so too does the number of alternating peaks and troughs of the inter-soliton potential energy. For soliton pairs with like sign, minima in $V(\Delta x)$ develop at $|\Delta x|\approx(I+1/2)\lambda_{\text{rot}}$, whereas for opposite-sign pairs they occur at $|\Delta x|\approx I\lambda_{\text{rot}}$, for $I=0,1,2,\dots$ . Correspondingly, maxima in $V(\Delta x)$ occur at $|\Delta x|\approx I\lambda_{\text{rot}}$ for like-sign solitons, and at $|\Delta x|\approx(I+1/2)\lambda_{\text{rot}}$ for opposite-sign solitons.
Since spin rotons are predicted across a broad range of dipolar mixtures [62], our arguments imply that multiple bound states are expected not only in dipolar-nondipolar mixtures but also more generally. Additional simulations in Appendix A corroborate this, showing such bound states for $\mu_{1}/\mu_{2}=-1$ and $\mu_{1}/\mu_{2}\approx 0.2$.
@@ III.6 Soliton collisions
True solitons can, in principle, emerge unscathed from collisions. However, the presence of dipolar interactions breaks integrability, so perfectly elastic collisions are not expected. Nevertheless, dark-antidark solitons may still survive collisions and thus serve as a probe of the inter-soliton potential.
Figure 5 shows four dark-antidark solitons in each panel—two slowly moving negative-core solitons and two stationary positive-core solitons—in (a) the nondipolar reference system and (b) a binary dipolar condensate. In the nondipolar case, like-sign solitons bounce off each other, whereas opposite-sign solitons pass through (as predicted in Ref. [38]). With dipolar interactions present, however, they always bounce, independent of relative sign. The bouncing arises from the periodic maxima in the inter-soliton potential $V$, which act as effective barriers. Crucially, universal low-velocity bouncing occurs because these maxima exist regardless of the soliton signs (see Sec. III.5). For like-sign solitons, these maxima occur at integer multiples of $\lambda_{\text{rot}}$; for opposite-sign solitons, half-integer multiples (see Fig. 4). Consequently, for the collisions shown in Fig. 5(b), the minimum separation before bouncing differs: $\Delta x\gtrsim\lambda_{\text{rot}}$ for like-sign collisions and $\Delta x\gtrsim 1.5\lambda_{\text{rot}}$ for opposite-sign collisions, suggesting that in both cases the second maxima are relevant at these velocities. These distinct collisional properties—especially when contrasted with the nondipolar case—could provide direct experimental evidence of the spin roton.
@@ IV Conclusion
Acknowledgements:— The authors acknowledge T. Billam, T. Bland, F. Ferlaino, L. B. Giacomelli, A. Madhusudan, N. Masalaeva, N. Parker, E. Poli, C. Qu and P. Senarath Yapa for insightful discussions. This research was funded by the Austrian Science Fund (FWF) [10.55776/P36850].
Appendix A Additional dipolar mixtures
To demonstrate the generality of our findings to the case where both $\mu_{1},\mu_{2}\neq 0$, we consider two additional dipolar configurations. In the first, we choose $\{\mu_{1},\mu_{2}\}=\{9.93,-9.93\}\mu_{\rm B}$ and an average per-component density $n_{0}=133~\text{atoms}~\mu\text{m}^{-1}$. Unless otherwise specified in the captions, the remaining parameters are the same as in Sec. II.3. Figure 6 shows the inter-soliton potential between opposite-sign solitons, with four minima present, each representing a different bound state. Figure 6(i) shows an example of one such bound state in an excited configuration, oscillating about the third minimum in Fig. 6. Importantly, both the inter-soliton potential and bound state behavior are qualitatively similar to the results presented in Fig. 3, despite the vastly different combination of dipole-dipole interaction parameters.
Figure 6: Inter-soliton potential versus soliton separation for a binary dipolar condensate with $\mu_{1}=-\mu_{2}$. Four minima are present ($x_{1}=0,x_{2}\approx 3.3,x_{3}\approx 6.7,x_{4}\approx 10.1$), corresponding to four regions where bound states can form. The inset shows example dynamics of the third bound state in an excited configuration, with the initial state position indicated as a cross in the main figure. Contact interaction parameters: $\{a_{11},a_{12},a_{22}\}/a_{0}=\{140,35,140\}$.
Figure 7: Example of a bound state for $\{\mu_{1},\mu_{2}\}=\{9.93,2\}\mu_{\rm B}$. This particular state becomes unbound at $t\omega\approx 70$. Contact interaction parameters: $\{a_{11},a_{12},a_{22}\}/a_{0}=\{140,121,140\}$.
In the second configuration, we set $\{\mu_{1},\mu_{2}\}=\{9.93,2\}\mu_{\rm B}$. Figure 7 shows an example of a bound state that remains stable until $t\omega\approx 70$. While these results may at first seem to suggest that a smaller difference between dipole moments may lead to shorter bound-state lifetimes, we also sometimes observed relatively short-lived bounds states for dipolar-nondipolar mixtures, hence further investigations are required to properly characterize lifetimes.
Appendix B Bogoliubov theory
@@
$\epsilon_{\pm}^{2}(k)=\frac{\epsilon_{1}^{2}+\epsilon_{2}^{2}}{2}\pm\frac{1}{2}\sqrt{(\epsilon_{1}^{2}-\epsilon_{2}^{2})^{2}+4G_{12}G_{12}\frac{\hbar^{4}k^{4}}{m^{2}}},$ (14)
@@
$\epsilon_{i}^{2}(k)=\frac{\hbar^{2}k^{2}}{2m}\left(\frac{\hbar^{2}k^{2}}{2m}+2G_{ii}(k)\right),$ (15)
@@
Appendix C Soliton effective mass
Here, we investigate the sign of the effective mass of dark-antidark soltions in binary dipolar BECs. In Fig. 8, we numerically calculate the energy $E$ (black crosses) of a single dark-antidark soltion as a function of its velocity $v$. For reference, we also plot the analyticaly predicted energy of a magnetic soliton in a binary nondipolar BEC (red line) [65]
$E_{\rm{ms}}(v)=\alpha\sqrt{1-v^{2}/c_{\rm{s}}^{2}},$ (17)
where $c_{\rm{s}}$ is the spin speed of sound and we choose $\alpha$ such that $E_{\rm{ms}}(v=0)=E(v=0)$. We can see that $E(v)$ decreases as $v$ increases, which indicates a negative effective mass according to the following expression, [72]
$m_{\rm{eff}}=\frac{1}{v_{\rm{s}}}\frac{dE}{dv_{\rm{s}}}.$ (18)
Figure 8: Energy vs velocity for a dark-antidark soliton in a dipolar-nondipolar mixture. Black crosses show the numerical predictions, while the red solid line shows the analytic prediction for a magnetic soliton in a binary nondipolar BEC. Contact interaction parameters: $\{a_{11},a_{12},a_{22}\}/a_{0}=\{140,95,140\}$.